How many pairs of (x,y) satisfy the equations 3x+6y=21 and 4x+8y=16

Grade:12th Pass

4 Answers

ABHISEK DASH
12 years ago

as 3/4=6/8  so these lines are parallel, i.e. have no common point, so 0 pairs of (x,y) will satisfy this

Jit Mitra
12 years ago

3x + 6y = 21

or, x+2y=7

 

4x+8y=16

x+2y=4

 

Inconsistent equations. No solution.

Rohan Das
11 years ago

dear suresh,

these two equations reaemble two parallel lines as a1/a2=b1/b2not equal to c1/c2.

so these two lines never intersect and thus have no solution

Akash Kumar Dutta
11 years ago

no solution since they are parallel lines
please post some tough questions next time
:)